A student weighs out a 1.91 g sample of NH4CH3COO , transfers it to a 100. mL volumetric flask, adds enough water to dissolve it and then adds water to the 100. mL tick mark.

What is the molarity of ammonium acetate in the resulting solution?



1
Expert's answer
2021-12-13T17:02:05-0500

Molar mass of Ammonium acetate "=77.0825g\/mol"


Number of moles in "100ml" solution "=\frac{1.91}{77.0825}"


Molarity "=\frac{1000}{100}×\frac{1.91}{77.0825}=0
.2478M"